Given the total cost function TC = 18q

2

+ 4q + 16,200

(a) Find marginal cost (MC) and average coat (AC) as functions of q

(b) Show that when MC < AC, AC is falling, and when MC > AC, AC is rising


1
Expert's answer
2021-03-17T08:31:27-0400

(a) The marginal cost funaction can be found as follows:


"MC=\dfrac{dTC}{dQ}=36q+4."

The average cost can be found as follows:


"AC=\dfrac{TC}{q}=18q+4+\dfrac{16200}{q}."

(b) If MC < AC, then AC is falling until MC = AC. Let's equate MC and AC:


"36q+4=18q+4+\dfrac{16200}{q},""18q^2-16200=0,""q=30."

If "q>30", MC>AC, therefore, AC is rising.